(Please use the same equations from question 1 for all 3 equations)

1) Suppose we have a supply curve with the equation y = x+1 and a demand curve with the equation y = -x+9. Using these two equation calculate the following and explain what these each mean: a) Market Equilibrium b) Consumer Surplus c) Producer Surplus

2) Suppose, due to a negative externality, that a quantity restriction has been placed on this market that restriction only 2 units of the product to be consumed/produced. Using the same supply/demand equations calculate the following: a) Change in Consumer Surplus b) Change in Producer Surplus c) Dead Weight Loss

3) Suppose now that a price ceiling allowing the product to be bought and sold at only $3/unit. Calculate the following a) Change in Consumer Surplus b) Change in Producer Surplus c) Dead Weight

PLEASE ANSWER ALL QUESTIONS
Answer #1

Suppose the supply cure ... and Y= x+1 the demand curve y = axta a market equilibrium requires Denang Demand = supply => xt =b) consues surplus is the demand care and abo Surplus is the area below the and above the price level MUL area of A.ABO SZx (Huce puce producers will pay is get is 43 and consumers will pay is $7. Thus, now, (a) the consues sur plus will be Area of Ay = $3 /unt in Now, Supply bstituting substituting y=- equation we get Y= x + 1 3 = x +1 => X = 2 in the demand and substitutHuce producn surplus falls by (8-2) 86 and Deadweight loss = Area of AROS - Lx (7-3) x (4-2) Iea Scanned with_ CamScanner